OpenStudy (anonymous):

1/128 times the definite integral from 0 to infinity of y^3 * e^(-y/8) Answer is 192. I get that we do integration by parts 3 times, but then I do not understand how we end up with 192.

OpenStudy (anonymous):

\[\int\limits_{0}^{\inf} = \lim_{t \rightarrow \infty}\int\limits_{0}^{t}\] lmk if this helps

OpenStudy (anonymous):

Thanks, I know how to do that general procedure, but I am having difficulty with this problem specifically

OpenStudy (anonymous):

The indefinite integral is: \[-8e^{-y/8}(y^3 + 24y^2 + 384y + 3072)\] thanks to wolfram. to take the limit as y->infinity, are you familiar with L'Hopital's Rule?

OpenStudy (anonymous):

Doesn't that just involve taking the derivative until it is simplified?

OpenStudy (anonymous):

If you could do it out for me, that would be great

OpenStudy (anonymous):

yes. just solved it and doesn't quite work

OpenStudy (anonymous):

actually it should. glad i pasted it in a notepad: pasting: it is for when the limit of f(x)/g(x) = 0/0 or infinity/infinity: \[\lim_{x -> \infty}\frac{ f(x) }{ g(x) } = \lim_{x -> \infty}\frac{ d/dx(f(x)) }{ d/dx(g(x)) }\] not to be confused with quotient rule. let's put it in the form f(x)/g(x) , limit is clearly infinity/infinity: \[-8 \lim_{y -> \infty} \frac{ y^3 + 24y^2 + 384y + 3072 }{ e^{y/8} }\] we derive numerator and denomninator respetively until it is no longer 0/0 or infty/infty ends up being:\[-8 \lim_{y -> \infty}\frac{ 6 }{ -\frac{ 1 }{ 512 } e^{y/8}} = 0\] so the entire infinity bound is 0. plugging in zero leaves you only with -(-8)*(3071)*(1/128) = 192

OpenStudy (anonymous):

3072** not 3071 lol

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

glad i could help ^_^

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.